# PWACG: Partial Wave Analysis Code Generator supporting Newton-conjugate gradient method

### v1

@@ 1 Introduction
In high-energy physics experiments, Partial Wave Analysis (PWA) serves as a vital data analysis method for extracting valuable information from measurements [1]. This technique involves analyzing the kinematic variables of final-state particles, particularly in multi-body processes, and integrating this information into the partial wave amplitudes for fitting [2]. It enables the determination of the presence of intermediate resonant states, along with their mass, width, parity, couplings, and relative phase angles between various amplitudes when considering interference. As a result, a large number of parameters conventionally need to be determined in a PWA [3], leading to a maximization problem in a high-dimensional parameter space. With the increasing scale of high-energy physics experimental data and the growing complexity of PWA models [4], PWA software faces long-term challenges in data processing capability and speed, thereby attracting sustained interest in its development.
@@
This challenge has been partially resolved by utilizing the continuous advancements in computer technology. Early versions of PWA programs were predominantly developed in C or Fortran language, focusing on CPU-based calculations, exemplified by the PWA2000 [5]. Constrained by the computing technology available at the time, these programs could only tackle decay problems involving small data sizes and relatively simple physical models [6]. In recent years, PWA programs utilizing C++ or Python as coding languages and based on Graphics Processing Units (GPUs) have emerged, such as GooFit [7] and GPUPWA [8]. GPUs offer a high degree of parallel processing capability, which, in comparison to traditional CPU programs, can significantly enhance computational speed and efficiency [9, 10], aligning well with the computational requirements of PWA. For instance, the GPUPWA package harnesses GPU parallel computing to expedite the calculation of likelihood functions and their derivatives, achieving a computational speed more than two orders of magnitude faster than traditional FORTRAN code [8]. However, with the rapid progress in particle accelerator and detector technologies, the volume of data generated by high-energy physics experiments has grown exponentially [11]. This surge in data has made it feasible to identify resonance states that were previously undetectable due to limited statistics, necessitating more complex PWA models to describe the data, entailing a significantly larger number of parameters [12]. The concurrent increase in the volume of experimental data and the number of parameters has placed greater demands on the performance of PWA software.
@@
Based on the experience in the field of artificial intelligence, it has been observed that the Newton conjugate gradient method [13, 14] can accelerate optimization by significantly reducing both computational load and memory usage [15], in comparison to the traditional standard Newton method or Quasi-Newton method [16]. Additionally, it has been noted that in the conjugate gradient calculation, automatic differentiation [17] is a valuable technique. Unlike the finite difference method [18], automatic differentiation can provide more precise results by minimizing rounding errors in floating-point operations, which is crucial in optimization when determining the descent direction [19]. However, mainstream PWA software still relies on the Minuit toolkit, including its FORTRAN version Minuit, C++ version Minuit2 [20], and Python version iMinuit [21]. The Minuit series primarily utilizes quasi-Newton methods for optimization [22], such as the BFGS [23] and DFP [24] methods. Even with the development of PWA software like TFPWA [25], which implements automatic differentiation, applying the Newton conjugate gradient method to practical PWA tasks remains highly challenging. To enhance PWA efficiency, a well-optimized package that combines the Newton conjugate gradient method and automatic differentiation is expected to significantly improve performance.
@@ 2.1 Architecture
To accurately derive the derivatives of the likelihood function, we turn to popular automatic differentiation tools currently in use. JAX [26], developed by Google, stands out as a high-performance machine learning library. It efficiently compiles and runs Python code using the XLA compiler, making it particularly well-suited for large-scale numerical computations. JAX takes full advantage of modern hardware, such as GPUs and TPUs, to accelerate computations, which is crucial for handling the complex calculations involved in partial wave analysis. Additionally, JAX supports ROCm and Apple’s deep learning chips, further broadening its applicability across different hardware platforms. Unlike PyTorch and TensorFlow, which are more focused on neural network computations, JAX can be seen as a GPU/TPU-accelerated version of NumPy. It offers a more fundamental approach in code implementation and superior automatic differentiation capabilities for mathematical functions, making it a natural choice for developing partial wave analysis software.
Throughout the partial wave analysis process, researchers need to continuously try adding or removing intermediate resonances to find the best partial wave analysis model that matches the data. Conventionally, we can introduce control parameters and use branches specific to the control parameters to determine whether to calculate a particular intermediate resonance. When using numerical differentiation to calculate derivatives, we only need to avoid calculating the derivatives of the control parameters. However, automatic differentiation software compiles the code before program execution, during which it deduces the differentiation calculation method for all variable parameters. The mathematical model corresponding to the branches for the control parameters is a step function, and the derivative corresponding to the step function diverges, which cannot be effectively handled by numerical methods. Therefore, automatic differentiation software does not support any type of branching command. To build an efficient partial wave analysis software, one feasible approach is to generate the corresponding partial wave analysis program based on the selected combination of resonances using code generation. Jinja2 is a popular Python template engine that provides powerful programming features such as variables, filters, and tags, making it highly suitable for handling complex logic and large templates. In the PWACG framework, we use jinja2 to combine various intermediate resonance templates into a complete partial wave analysis model based on the configuration file provided by the user.
@@ 2.2 Implementation and Performance Optimization of Computational Templates
[…]
@@ 2.3 Multi-GPU Support
$F_{m}=\sum_{i=1}^{N_{Data,m}}\ln\left(\sum_{\mu=1}^{2}A^{\mu}(\xi_{i})A^{*\mu}(\xi_{i})\right),\leavevmode\nobreak\ G_{m}=\sum_{j=1}^{N_{MC,m}}\sum_{\mu=1}^{2}A^{\mu}(\xi_{j})A^{*\mu}(\xi_{j})$ (5)
@@ 3 Performance Test
PWACG generates high-performance PWA code that can rapidly and accurately compute the likelihood, first-order derivative, and HVP of any PWA model. This makes it suitable for various optimization methods. To comprehensively evaluate the performance of PWACG, we constructed a PWA model to describe the decay process $\psi(3686)\rightarrow\phi K^{+}K^{-}$ [27]. This model encompasses intermediate processes such as $\psi(3686)\to\phi f$, $f\to K^{+}K^{-}$, where $f$ represents various scalar or tensor resonances including $f_{0}(980)$ [28, 29], $f_{2}(1270)$ [30], $f_{2}^{\prime}(1525)$ [31], $f_{0}(1710)$ [32], and $f_{2}(2150)$ [33]. The fraction, mass, and width of these resonances are detailed in Table 1, while the phase angles and couplings of individual partial wave amplitudes are not included here for brevity. A Monte Carlo (MC) sample comprising 10,000 events based on this physical model was generated. The corresponding Dalitz plot and $K^{+}K^{-}$ invariant mass spectrum are depicted in Fig. 1 and Fig. 2, respectively. This model encompasses both broad and narrow resonances, with relatively strong interference between the resonances, presenting a challenging scenario for PWA. It is important to note that this MC sample set is utilized for meticulously testing the performance of the PWA program, rather than providing a complete representation of the actual three-body decay $\psi(2S)\rightarrow\phi K^{+}K^{-}$.
@@
  • iMinuit+FDM: Calculate derivatives using finite difference method, and use iMinuit for optimization.
@@
We randomly generated 300 sets of parameters and used them as the fitting starting points with the three methods. The distributions of the resulting $-\ln L$ values are shown in Fig. 4. On average, iMinuit+AutoDiff finishes the fitting the fastest, followed by iMinuit+FDM, due to potential precision loss near the fitting endpoint with the finite difference method, resulting in a slower convergence rate. NTCG+AutoDiff, while slower to finish, exhibits a higher success rate in reaching the optimal point compared to the other two methods. The global optimum value of $-\ln L$ is $-4110$, with results below $-4109$ considered the global optimum for statistical purposes. The success rate of NTCG+AutoDiff in finding the global optimum is 35.7%, while iMinuit+AutoDiff and iMinuit+FDM have success rates of only 6.3% and 4.3%, respectively, as shown in Fig. 5.
Given that PWA is a non-convex optimization problem, finding the global optimum requires attempting fits from randomly initialized parameters, and the success rate of each attempt significantly impacts the analysis efficiency. Assuming the success rate of each attempt is $\alpha$, then after N attempts, the probability of missing the global optimum is $(1-\alpha)^{N}$. Therefore, to ensure the probability of missing the global optimum is less than 1/100, NTCG+AutoDiff needs 11 attempts, iMinuit+AutoDiff needs 71 attempts, and iMinuit+FDM needs 105 attempts. Although NTCG+AutoDiff takes longer for a single fit, its substantially higher probability of finding the global optimum in each fit makes it highly competitive overall.
In more complex PWA models with a larger parameter space, the probability of finding the global optimum through random sampling would be lower, further emphasizing the advantages of the NTCG+AutoDiff method. Additionally, for NTCG+AutoDiff, an “early stopping [34]” strategy similar to deep learning optimization, where the fitting process is terminated if it exceeds 500 seconds, can be adopted, minimally impacting the success rate of finding the global optimum while significantly saving time.
Figure 4: Scatter plot of $-\ln L$ versus time with the results of 300 fits for each optimization method.
Figure 5: Fitting times needed for three different methods, based on 300 fits per method.
@@
Multi-GPU parallel computing can effectively mitigate the limitations of insufficient memory on a single GPU and substantially enhance fitting efficiency. In order to fully showcase the acceleration capabilities of multi-GPU parallel computing, we conducted parallel tests using different numbers of GPUs while maintaining the model and data scale unchanged. It is important to note that, at present, PWACG exclusively supports multi-GPU computation on a single node. However, in the future, support for multi-GPU parallelism across nodes could be implemented through MPI. Figure 7 illustrates the distribution of time spent on fitting using different numbers of GPUs under the condition of 1,200,000 integration instances. The results demonstrate that PWACG has undergone significant optimizations for multi-GPU parallel computing, with multiple GPUs delivering substantial performance improvements.
@@
Figure 7: Time consumed for parallel fitting on multiple GPUs.

### v2

@@ Abstract
Program Summary
  • Program Title: PWACG: Partial Wave Analysis Code Generator
  • Licensing provisions: This software is distributed under the MIT License.
  • Programming language: Python
  • External routines/libraries: jaxlib, jax, jinja2, matplotlib, numpy, scipy
  • Nature of problem: The program addresses the need for high-performance computational tools in high-energy physics partial wave analysis (PWA). It introduces the Newton-conjugate gradient method for optimization, enhancing the accuracy and stability of fits.
  • Solution method: PWACG employs code generation and automatic differentiation techniques to automate the creation of PWA code. It leverages the computational capabilities of JAX for efficient execution and supports advanced features such as multi-GPU computation.
  • GitHub repository: https://github.com/caihao/PWACG
@@ 1 Introduction
In high-energy physics experiments, Partial Wave Analysis (PWA) serves as a vital data analysis method for extracting valuable information from measurements [peters2006primer]. This technique involves analyzing the kinematic variables of final-state particles, particularly in multi-body processes, and integrating this information into the partial wave amplitudes for fitting [klempt2017partial]. It enables the determination of the presence of intermediate resonant states, along with their mass, width, parity, couplings, and relative phase angles between various amplitudes when considering interference. As a result, a large number of parameters conventionally need to be determined in a PWA [Krinner2018Ambiguities], leading to a maximization problem in a high-dimensional parameter space. With the increasing scale of high-energy physics experimental data and the growing complexity of PWA models [yuan2019besiii], PWA software faces long-term challenges in data processing capability and speed, thereby attracting sustained interest in its development.
@@
This challenge has been partially resolved by utilizing the continuous advancements in computer technology. Early versions of PWA programs were predominantly developed in C or Fortran language, focusing on CPU-based calculations, exemplified by the PWA2000 [cummings2003object]. Constrained by the computing technology available at the time, these programs could only tackle decay problems involving small data sizes and relatively simple physical models [aubert2005amplitude]. In recent years, PWA programs utilizing C++ or Python as coding languages and based on Graphics Processing Units (GPUs) have emerged, such as GooFit [Andreassen_2014] and GPUPWA [Berger2010Partial]. GPUs offer a high degree of parallel processing capability, which, in comparison to traditional CPU programs, can significantly enhance computational speed and efficiency [Park2009An, Owens2008GPU], aligning well with the computational requirements of PWA. For instance, the GPUPWA package harnesses GPU parallel computing to expedite the calculation of likelihood functions and their derivatives, achieving a computational speed more than two orders of magnitude faster than traditional FORTRAN code [Berger2010Partial]. However, with the rapid progress in particle accelerator and detector technologies, the volume of data generated by high-energy physics experiments has grown exponentially [Apostolakis2022Detector]. This surge in data has made it feasible to identify resonance states that were previously undetectable due to limited statistics, necessitating more complex PWA models to describe the data, entailing a significantly larger number of parameters [FCpwa]. The concurrent increase in the volume of experimental data and the number of parameters has placed greater demands on the performance of PWA software. Recent advancements also include the development of frameworks such as AmpliTF [amplitf_github] and ComPWA [compwa_docs], which leverage TensorFlow and JAX, respectively, to enhance the efficiency and flexibility of PWA tasks, paving the way for more robust fitting and modeling tools.
@@
Based on the experience in the field of artificial intelligence, it has been observed that the Newton conjugate gradient method [Yang2009Newton-conjugate-gradient, Sherali1990Conjugate] can accelerate optimization by significantly reducing both computational load and memory usage [song2022modeling], in comparison to the traditional standard Newton method. Moreover, while Quasi-Newton methods [Nocedal2021Quasi-Newton] approximate the Hessian matrix using gradient information, the Newton conjugate gradient method benefits from exact second-order information, leading to higher efficiency in locating global optima. Additionally, it has been noted that in the conjugate gradient calculation, automatic differentiation [baydin2018automatic] is a valuable technique. Unlike the finite difference method [Mazumder2016The], automatic differentiation can provide more precise results by minimizing rounding errors in floating-point operations, which is crucial in optimization when determining the descent direction [Ruder2016An]. However, mainstream PWA software still relies on the Minuit toolkit, including its FORTRAN version Minuit, C++ version Minuit2 [Hatlo2004Developments], and Python version iMinuit [iminuit_github]. The Minuit series primarily utilizes quasi-Newton methods for optimization [James1984MINUIT-a], such as the BFGS [Badem2018A] and DFP [Avriel1976] methods. Even with the development of PWA software like TFPWA [tfpwa_github]and zfit [eschle2019zfit], which implements automatic differentiation, applying the Newton conjugate gradient method to practical PWA tasks remains highly challenging. To enhance PWA efficiency, a well-optimized package that combines the Newton conjugate gradient method and automatic differentiation is expected to significantly improve performance.
@@ 2.1 Architecture
To accurately derive the derivatives of the likelihood function, we turn to popular automatic differentiation tools currently in use. JAX [jax2018github], developed by Google, stands out as a high-performance machine learning library. It efficiently compiles and runs Python code using the XLA compiler, making it particularly well-suited for large-scale numerical computations. JAX takes full advantage of modern hardware, such as GPUs and TPUs, to accelerate computations, which is crucial for handling the complex calculations involved in partial wave analysis. Additionally, JAX supports ROCm and Apple’s deep learning chips, further broadening its applicability across different hardware platforms. Unlike PyTorch and TensorFlow, which are more focused on neural network computations, JAX can be seen as a GPU/TPU-accelerated version of NumPy. It offers a more fundamental approach in code implementation and superior automatic differentiation capabilities for mathematical functions, making it a natural choice for developing partial wave analysis software.
Throughout the partial wave analysis process, researchers need to continuously try adding or removing intermediate resonances to find the best partial wave analysis model that matches the data. Conventionally, we can introduce control parameters and use branches specific to the control parameters to determine whether to calculate a particular intermediate resonance. When using numerical differentiation to calculate derivatives, we only need to avoid calculating the derivatives of the control parameters. Nonetheless, to attain high computational performance, our implementation necessarily relies on JAX’s just-in-time (JIT) compilation mechanism. JIT compilation works by tracing the execution of a Python function and transforming it into a static computation graph, which is then optimized and executed by the XLA (Accelerated Linear Algebra) compiler. A critical limitation of this static-graph-based approach is its inability to support data-dependent control flow, particularly when expressed using standard Python conditional statements. This is because such constructs introduce dynamic structural changes to the computation graph at runtime. As a result, this constraint fundamentally conflicts with XLA’s requirement for static computation graphs, posing a significant engineering challenge in designing flexible and efficient algorithms. To build an efficient partial wave analysis software, one feasible approach is to generate the corresponding partial wave analysis program based on the selected combination of resonances using code generation. Jinja2 is a popular Python template engine that provides powerful programming features such as variables, filters, and tags, making it highly suitable for handling complex logic and large templates. In the PWACG framework, we use jinja2 to combine various intermediate resonance templates into a complete partial wave analysis model based on the configuration file provided by the user.
@@ 2.2 Implementation and Performance Optimization of Computational Templates
$L=\prod_{i=1}^{N_{Data}}\frac{\omega(\xi_{i},\alpha)}{\int d\xi\omega(\xi,\alpha)\epsilon(\xi)}$ (2)
@@ 2.3 Multi-GPU Support
$F_{m}=\sum_{i=1}^{N_{Data,m}}\ln\left(\sum_{\mu=1}^{2}A^{\mu}(\xi_{i})A^{*\mu}(\xi_{i})\right),\penalty 10000\ G_{m}=\sum_{j=1}^{N_{MC,m}}\sum_{\mu=1}^{2}A^{\mu}(\xi_{j})A^{*\mu}(\xi_{j})$ (5)
@@ 3 Performance Test
PWACG generates high-performance PWA code that can rapidly and accurately compute the likelihood, first-order derivative, and HVP of any PWA model. This makes it suitable for various optimization methods. To comprehensively evaluate the performance of PWACG, we constructed a PWA model to describe the decay process $\psi(3686)\rightarrow\phi K^{+}K^{-}$ [ZouBS]. This model encompasses intermediate processes such as $\psi(3686)\to\phi f$, $f\to K^{+}K^{-}$, where $f$ represents various scalar or tensor resonances including $f_{0}(980)$ [flatte1, flatte2], $f_{2}(1270)$ [Shchegelsky:2006et], $f_{2}^{\prime}(1525)$ [Longacre:1986fh], $f_{0}(1710)$ [Belle:2013eck], and $f_{2}(2150)$ [WA102:2000lao]. The fraction, mass, and width of these resonances are detailed in Table 1, while the phase angles and couplings of individual partial wave amplitudes are not included here for brevity. A Monte Carlo (MC) sample comprising 10,000 events based on this physical model was generated. The corresponding Dalitz plot and $K^{+}K^{-}$ invariant mass spectrum are depicted in Fig. 1 and Fig. 2, respectively. This model encompasses both broad and narrow resonances, with relatively strong interference between the resonances, presenting a challenging scenario for PWA. It is important to note that this MC sample set is utilized for meticulously testing the performance of the PWA program, rather than providing a complete representation of the actual three-body decay $\psi(2S)\rightarrow\phi K^{+}K^{-}$.
@@
  • iMinuit+FDM: Calculate derivatives using finite difference method(FDM), and use iMinuit for optimization.
@@
We randomly generated 300 sets of parameters and used them as the fitting starting points with the three methods. A fundamental criterion for a valid performance comparison of these optimization methods is to evaluate them under an equivalent convergence precision. In our example PWA model, our empirical tests reveal significant differences in the maximum achievable precision for each approach:
  • iMinuit+FDM: The convergence of the function to be minimized (FCN) is limited to a precision of approximately $10^{-2}$. Setting a smaller value leads to convergence difficulties and increases the computation time, without yielding any improvement in precision. In iMinuit, the parameter toltol can be adjusted to affect the max estimated distance to minimum (EDM), thereby controlling the convergence precision [iminuit_web].
  • iMinuit+AutoDiff: Since AutoDiff provides more accurate derivatives, this method can achieve an FCN convergence precision up to $10^{-5}$.
  • NTCG+AutoDiff: In the NTCG fitting method provided by scipy.optimize, we influence the convergence precision of FCN by controlling the convergence precision of fitting parameters xtolxtol [SciPyOptimize]. The default value of xtolxtol is $10^{-8}$, and we can relax xtolxtol restriction to provide a more reasonable comparison condition with the two methods above.
In Figure 4, we compare the fitting capabilities of these three methods. Here we set the xtolxtol parameter of NTCG+AutoDiff to $10^{-5}$ to achieve an FCN convergence precision close to that of iMinuit+AutoDiff. The distributions of the resulting $\ln L$ values are shown in Fig. 4(a). The global optimum value of $\ln L$ is 4110-4110, with results below 4109-4109 considered the global optimum for statistical purposes. Among these methods, NTCG+AutoDiff has the highest success rate in finding the global optimum at 6.67%, followed by iMinuit+AutoDiff at 1.33%, while iMinuit+FDM has a rate of only 0.33%, as shown in Fig. 4(b).
(a) Scatter plot of $\ln L$ versus time.
(b) $\ln L$ distributions of 300 fits.
Figure 4: The results of 300 fits for three different optimization methods. The convergence precision of FCN for method iMinuit+FDM was set to $10^{-2}$ , while the methods iMinuit+AutoDiff and NTCG+AutoDiff was set to $10^{-5}$.
In actural PWA probelm, NTCG+AutoDiff allows us to set a much lower xtolxtol, below $10^{-8}$. This setting effectively helps in avoiding local minimal, ultimately achieving a success rate of 35%, as shown in Fig. 5(a) and Fig. 5(b) . Additionally, for NTCG+AutoDiff, an “early stopping [dossa2021empirical]” strategy similar to deep learning optimization, where the fitting process is terminated if it exceeds 500 seconds, can be adopted, minimally impacting the success rate of finding the global optimum while significantly saving time.
(a) Scatter plot of $\ln L$ versus time.
(b) $\ln L$ distributions of 300 fits.
Figure 5: Performance comparison of NTCG+AutoDiff at $xtol=10^{-8}$ and $xtol=10^{-5}$
@@
Multi-GPU parallel computing can effectively mitigate the limitations of insufficient memory on a single GPU and substantially enhance fitting efficiency. In order to fully showcase the acceleration capabilities of multi-GPU parallel computing, we conducted parallel tests using different numbers of GPUs while maintaining the model and data scale unchanged. It is important to note that, at present, PWACG exclusively supports multi-GPU computation on a single node. However, in the future, support for multi-GPU parallelism across nodes could be implemented through MPI. Fig. 7 illustrates the average time consumed in different numbers of GPUs under the condition of 1,200,000 integration instances. The results demonstrate that PWACG has undergone significant optimizations for multi-GPU parallel computing, with multiple GPUs delivering substantial performance improvements.
@@
Figure 7: Average time consumed for parallel fitting under different GPU configurations.